OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -2 and 16, respectively. a_n = 1 • (-2)n - 1 a_n = 1 • 2n a_n = 1 • (-2)n + 1 a_n = 1 • 2n - 1

OpenStudy (aum):

nth term of a geometric sequence is: \(\large a_n = ar^{n-1}\) ----- (1) \(\large a_2 = ar^{2-1} = ar = -2\) ----- (2) \(\large a_5 = ar^{5-1} = ar^4 = 16\) ----- (3) Divide (3) by (2) and solve for 'r': Then put 'r' in (2) and solve for 'a' Put 'a' and 'r' in (1).